ganeshie8 (ganeshie8):

Find any 3 consecutive integers such that each of them is divisible by a perfect square > 1

ganeshie8 (ganeshie8):

For example : first integer can be divisible by 4, second integer by 25 and third integer by 49

OpenStudy (anonymous):

hmm,if we have an integer like \(t\) then \( t \equiv 0,1,2,3 ~ (mod ~ 4) \) so \( t^2 \equiv 0,1 \) but as we have 3 consecutive numbers none of the numbers can be divisible by the same numbers : suppose \( s^2 | a \) and \( s^2 | a+2\) so \( s^2 | 2 \) and it can never happen.

OpenStudy (anonymous):

therefore we have 3 perfect square which all of them are different and we have 2 cases for being congruent (mod 4) , using pigeonhole principle we can say two of the square numbers are congruent (mod 4)

OpenStudy (anonymous):

well,maybe this can be used as the idea ;D work on it.

OpenStudy (anonymous):

yes! we can't say that two perfect squares are divisble by 4 because then we can say 4|a and 4|a+1 or a+2 and it can't happen so, \( t^2 \equiv s^2 \equiv 1 ~ (mod ~ 4 ~ )\)

ganeshie8 (ganeshie8):

thats right! the divisors must be distinct as the same number cannot divide a, a+1 and a+2

OpenStudy (anonymous):

well @ganeshie8 i'm not sure this can be a good idea or not!

ganeshie8 (ganeshie8):

I am thinking something like this suppose 4 divides a 9 divides a+1 25 divides a+2

OpenStudy (anonymous):

u only need one set of numbers?

ganeshie8 (ganeshie8):

yes just looking for one set of solution :)

OpenStudy (mathmath333):

\(\Large \begin{align} \color{black}{ 548 ~mod ~2^2 \equiv 0\hspace{.33em}\\~\\ 549~ mod ~3^2 \equiv 0\hspace{.33em}\\~\\ 550 ~mod ~5^2 \equiv 0\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

correct ;) how did u find?

ganeshie8 (ganeshie8):

thats very fast ! xD

OpenStudy (mathmath333):

yes , actually their can be many of examples of these this is an example of chinese remainder theorm yesterday @ganeshie8 taught me a problem these so i had the feeling

OpenStudy (anonymous):

oh! yes chinese remiander thm but @mathmath333 that thm only say that there most be an integer which is divisble by n perfect square.

OpenStudy (anonymous):

it won't give u the numbers...

OpenStudy (anonymous):

* must

OpenStudy (mathmath333):

well u could consider \(n^2\) as any integers ,ex,4 9 ,36, 49,64,81

OpenStudy (mathmath333):

another example \(\Large \begin{align} \color{black}{ 15849 ~mod ~3^2 \equiv 0\hspace{.33em}\\~\\ 15850 ~mod ~5^2 \equiv 0\hspace{.33em}\\~\\ 15851~ mod ~11^2 \equiv 0\hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

Looks neat :) here is one way to work it with out using chinese remainder thm : \[\begin{align} a&\equiv 0 \pmod 4\\~\\ a+1&\equiv 0 \pmod{9}\\~\\ a+2&\equiv 0 \pmod{25} \end{align}\] from the first congruence we have \(a = 4m\) plugging that in second congruence we get \(4m + 1\equiv 0 \pmod{9}\) solving gives \(m = 2 + 9n\) plugging this back into \(a\) we get \(a = 4(2+9n)\) we can plug this into third congruence and solve...

OpenStudy (mathmath333):

yes^^

OpenStudy (anonymous):

@ganeshie8 thanks ;) @mathmath333 nice way ;)

OpenStudy (mathmath333):

i think the value of the three squares should be coprime or should be squares of prime numbers ,is that correct @ganeshie8

ganeshie8 (ganeshie8):

thats right! to apply chinese remainder theorem we need the moduli to be corprime to each other

OpenStudy (mathmath333):

yes so something like \(2^2,4^2,16^2\) wont work here

ganeshie8 (ganeshie8):

\(a\equiv 0 \pmod {2^2} \) \(a+1\equiv 0 \pmod {4^2} \) itself is not solvabel because the gcd of 2^2 and 4^2 is 2^2 and 2^2 does not divide -1

ganeshie8 (ganeshie8):

we have this rule : \[x\equiv a \pmod n\\x\equiv b \pmod m\] this system is solvable only when \(\gcd(m,n)\) divides \(a-b\)

ganeshie8 (ganeshie8):

I'll post a separate question for its proof it is fun :)

OpenStudy (anonymous):

from the first equation we have \( x = nk + a \) and from the second \( x = mk^* + b\) so \(a-b + nk - mk^*= 0 \) therefore \( a-b = mk^* - nk\) but \( gcd(m,n) | mk^* \) and \(gcd(m,n) | nk \) so \( a-b|gcd (m,n) \)

OpenStudy (anonymous):

but besides this we have to prove that when \( a-b | gcd(m,n)\) then that system is solvable.

OpenStudy (anonymous):

then we can say that system is resolvable when and only when \( a-b|gcd(m,n) \)

ganeshie8 (ganeshie8):

thats exactly same as the proof im thinking of ! xD \[a-b = mk^* - nk\] suppose gcd(m,n) = d, then we can factor this out from the right hand side : \[a-b = d(stuff)\] clearly \(a-b\) is a multiple of \(d\) which is same as saying \(d\) divides \(a-b\)

OpenStudy (anonymous):

i'm thinking for the second part of the proof.suppose a-b|d then prove that system is resolvable.

ganeshie8 (ganeshie8):

Ahh yes that looks bit involved

OpenStudy (anonymous):

yes,maybe doing the same thing as we do for proving chineese remiander thm!

OpenStudy (mathmath333):

@PFEH.1999 \(\Large \begin{align} \color{black}{ a~ mod ~2^2 \equiv 0\hspace{.33em}\\~\\ a+1 ~mod ~3^2 \equiv 0\hspace{.33em}\\~\\ a+2 ~mod ~5^2 \equiv 0\hspace{.33em}\\~\\ }\end{align}\) this is same as \(\Large \begin{align} \color{black}{ a ~mod ~2^2 \equiv 0\hspace{.33em}\\~\\ a ~mod ~3^2 \equiv 8\hspace{.33em}\\~\\ a~ mod ~5^2 \equiv 23\hspace{.33em}\\~\\ }\end{align}\) thats why we apply chinese rem thrm

OpenStudy (anonymous):

and from this we can find a ?

OpenStudy (mathmath333):

yes of cos

OpenStudy (anonymous):

yes thank toy ;) beautiful way ;)

OpenStudy (mathmath333):

oh lol sry it \(\huge 548\) please check it

OpenStudy (anonymous):

correct

OpenStudy (mathmath333):

thnks

OpenStudy (anonymous):

and would u write ur complete solution i don't know why i get wrong answer!

OpenStudy (mathmath333):

yes here is a problem like that and their are 3 to 4 methods for solving it, have a look http://openstudy.com/study#/updates/54aabff5e4b0b83d9dca0684

OpenStudy (anonymous):

thanks ;)

OpenStudy (anonymous):

thinking of Quadratic residue

OpenStudy (anonymous):

\(n|p_1^2\\n-1|p_2^2\\n+1|p_3^2\\n(n-1)(n+1)|(p_1p_2p_3)^2~~~~~\text{s.t p1,p2,p3 can be different or the same }\)

OpenStudy (anonymous):

@mathmath333 , a=4k , 4k = 8 ( mod 9) so k=2 (mod 9) so k=9s + 2 and a would be 36s + 8 thus 25 | 36s - 15 so 25|12s - 5

OpenStudy (anonymous):

how to solve s and plug in to find a?

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ a ~mod ~2^2 \equiv 0\hspace{.33em}\\~\\ a ~mod ~3^2 \equiv 8\hspace{.33em}\\~\\ a ~mod~ 5^2 \equiv 23\hspace{.33em}\\~\\ \normalsize \text{by the chinese remainder theorm}\hspace{.33em}\\~\\ a=x+y+z~(mod~m)\hspace{.33em}\\~\\ x=0\cdot 9\cdot 25\left( (9\cdot 25)^{-1}~(mod~4) \right)\hspace{.33em}\\~\\ x=0\hspace{.33em}\\~\\ y=8\cdot 4\cdot 25\left( (4\cdot 25)^{-1}~(mod~9) \right)\hspace{.33em}\\~\\ y=8\cdot 4\cdot 25\left(1 \right)\hspace{.33em}\\~\\ z=23\cdot 4\cdot 9\left( (4\cdot 9)^{-1}~(mod~25) \right)\hspace{.33em}\\~\\ z=23\cdot 4\cdot 9\left( 16 \right)\hspace{.33em}\\~\\ m= 4\cdot 9\cdot 25\hspace{.33em}\\~\\ a=\left(0+8\cdot 4\cdot 25\left(1 \right)+23\cdot 4\cdot 9\left( 16 \right)\right) ~(mod~4\cdot 9\cdot 25)\hspace{.33em}\\~\\ a=\dfrac{8\cdot 4\cdot 25 }{4\cdot 9\cdot 25}+\dfrac{23\cdot 4\cdot 9\left( 16 \right) }{4\cdot 9\cdot 25}\hspace{.33em}\\~\\ a=\dfrac{8 }{ 9}+\dfrac{23\left( 16 \right) }{ 25}\hspace{.33em}\\~\\ a=\dfrac{-1 }{ 9}+\dfrac{-7}{ 25}\hspace{.33em}\\~\\ a=\dfrac{-88 }{ 9\cdot 25}\\~\\ a=\dfrac{-88\cdot 4 }{ 9\cdot 25\cdot 4 }\\~\\ a=\dfrac{-352}{ 900 }\\~\\ a=\dfrac{900-352}{ 900 }\\~\\ a=\dfrac{548}{ 900 }\\~\\ a=548 }\end{align}\)